# Differentially Private Matchings

### v1

@@ 1.1.4 Continual Release
We give results for two different types of streams in the continual release model. The first type of stream we consider is the arbitrary edge-order stream where the edges in the stream appear one by one in arbitrary order. In particular, in arbitrary edge-order streams, we show the following results for edge- and node-DP. In our results below, we give bicriteria approximate solutions where an $(\psi,\phi)$-approximation is one with a $\psi$-multiplicative error and $\phi$-additive error. Our approximations hold, with high probability, for every update.
@@
Let $\eta\in(0,1]$ and $\varepsilon\in(0,1)$. There is an $\varepsilon$-edge DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
@@
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log^{3}(n)}{\eta^{2}\varepsilon}\right)$and (ii) each implicit solution contains a
$\left(2+\eta,O\left(\frac{\alpha\log^{3}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
The second type of stream we consider is the arbitrary adjacency-list order model [kallaugher2019complexity, goldreich2009proximity, czumaj2016relating, mcgregor2016better] in which nodes arrive in arbitrary order and once a node $v$ arrives as an update in the stream all edges adjacent to $v$ arrive in an arbitrary order as edge updates immediately after the arrival of $v$. In this model, we give the following results.
@@
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, there is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs, with high probability, an implicit $b$-matching containing a maximal matching with additive error of at most $1$.
Corollary 1.12.
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, there is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs, with high probability, an implicit $b$-match containing a $(2,1)$-approximate maximum matching.
@@
Model $b$ Approximation Solution Type Bound Type
$(\varepsilon,\delta)$-edge DP $O(1)$ $\Omega(n/b)$ Explicit Lower Bound
$\varepsilon$-edge DP $\Omega(\log(n)/\varepsilon)$ $(2+\eta)$ Implicit Lower Bound
$\varepsilon$-LEDP ($n$ rounds) $O(\log(n)/\varepsilon)$ $2$ Implicit Upper Bound
$\varepsilon$-LEDP ($O(\log n)$ rounds) $O(\log^{2}(n)/\varepsilon)$ $2$ Implicit Upper Bound
$\varepsilon$-node DP $O\left(\frac{\alpha\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$ Implicit Upper Bound
$\varepsilon$-node DP (bipartite) $s=\Omega(\log(n)/\varepsilon)$ $(1+\eta)$, $s$-matching Implicit Upper Bound
Edge-Order $\varepsilon$-Edge DP $O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Implicit Upper Bound
@@
Edge-Order $\varepsilon$-Node DP $O\left(\frac{\alpha\log^{3}(n)}{\eta^{2}\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\alpha\log^{3}(n)}{\eta\varepsilon}\right)\right)$ Implicit Upper Bound
@@
Adj-List Edge $O(\log(n)/\varepsilon)$ $(2,1)$ Implicit Upper Bound
@@
Table 1: Summary of results on differentially private $b$-matching. All approximations are given in terms of the optimum maximum matching in the input graph.
@@ 1.2.1 Edge Privacy
Our main algorithms (Theorem 1.2 and Theorem 1.4) take inspiration from the well-known greedy algorithm as well as distributed algorithms for maximal matching. In particular, the traditional greedy algorithm for maximal matching iterates through the nodes in the graph in an arbitrary order and matches the nodes greedily whenever there exists an unmatched node in the current iteration. However, this algorithm is clearly not private because each matched edge reveals the existence of a true edge in the graph. Thus, we must have some way of releasing public subsets of vertices that include both true edges and non-edges.
The algorithm from Theorem 1.2 proceeds in $n$ rounds using an arbitrary ordering of the nodes in the graph. We give as input a threshold $b$ which is a threshold for the number of edges in the matching each vertex is incident to. In the order provided by the ordering, each node $v$ privately checks using the Multidimensional AboveThreshold (MAT) technique introduced in [DLL23] whether it has reached its number of matched neighbors threshold. If it has not, then it asks all of its neighbors ordered after it whether they will match with it. Each of these neighbors first determines using MAT whether they have exceeded their threshold for matching and then flips a coin (using progressively exponentially decreasing probability) to determine whether to be part of the set to match with $v$. All of these procedures can be done locally privately in the LEDP model. Our privacy proof depends on recent work on concurrent composition [vadhan2021concurrent], and to the best of our knowledge is the first use of this technique in private graph algorithms.
@@
To speed up our algorithm to run in polylogarithmic rounds (Theorem 1.4), we need to allow many vertices to simultaneously propose matches to their neighbors (rather than just one at a time). Intuitively, this can be done in ways similar to classical parallel and distributed algorithms for related problems like Boruvka’s parallel MST algorithm [Boruvka]. In these algorithms, nodes randomly choose whether to propose or to listen. Then, proposer nodes send proposals, receivers decide which of their proposals to accept, and receivers communicate the acceptances back to the proposers (see [DHIN19] for a recent example of this in distributed settings). However, the added LEDP constraint makes this process significantly more difficult. Note, for example, that even communicating “with neighbors” violates privacy, since the transcript would then reveal who the neighbors are! Hence greater care must be taken regarding the messages each node transmits. Fortunately, by using techniques gleaned from our $n$-round algorithm, ideas from basic parallel algorithms, and improved analyses, our implementation goes through.
The crux of our $O(\log n)$ round distributed algorithm relies on our Public Vertex Subset Mechanism. We use this mechanism to release proposal sets on the part of the proposers and match sets on the part of the receivers. The proposers release public proposal sets that contain proposals to its neighbors to match. Using the Public Vertex Subset Mechanism, receivers can privately discern which nodes have proposed. Then, receivers also publish a public subset of nodes as their match sets. Using this mechanism, the proposers can then (privately) recognize which receivers accepted their proposals. Performing multiple rounds of this procedure leads to our desired Theorem 1.4. Our novel utility analyses depends on a noisy version of a charging argument based on edge orientation implicitly described in [khuller1994primal].
@@ 1.2.2 Node Privacy
So if we could compute and release a differentially private matching sparsifier, we would be done. Unfortunately, we show in Section 9.2 that this is impossible, by providing a strong lower bound against private matching sparsification (Theorem 9.2). However, we show that we can use these sparsifiers in a node-DP algorithm even though releasing them does not satisfy DP! Intuitively this is because given two node-neighboring graphs, computing a sparsifier of them can bring the distributions of the outputs “closer together”. So we cannot release such a sparsifier, but it does not actually violate privacy as long as we do not release it. In particular, although these sparsifiers cannot be released publicly, they decrease the “edge-distance” between node-neighboring graphs, where the edge-distance is the number of edges that differ between neighboring graphs.
Bipartite Graphs. (Section 8)
The starting point of our algorithm for node-private bipartite matchings is the same deferred acceptance type algorithm which [hsu2014private] base their private algorithm on. In their private implementation of the algorithm, they use continual counters [dwork2009complexity, CSS11] to keep track of the number of bidders each item is matched to. This is necessary for their implementation since their implicit solutions need to know the number of bidders an item is matched with at every iteration of the algorithm. We show that in a different private implementation of the algorithm, the counts of the number of bidders an item is matched to is only needed $O(1/\eta^{2})$ different times (where $\eta\in(0,1]$ is present in the $(1+\eta)$-approximation factor). Furthermore, a clever use of MAT to iteratively check if this count exceeds the threshold $s$ (where $s$ is the matching factor) suffices for the implementation, enabling us to substantially improve the error of the counts used. As before, our privacy analysis makes use of the recent work on concurrent composition theorems [vadhan2021concurrent].
1.2.3 Continual Release (Section 7)
@@ 1.3 Related Work
Differentially private estimations of the size of the maximum matching have been recently studied in various privacy models including the continual release model [FHO21, wagaman2024time, raskhodnikova2024fully] and the sublinear model [BGM22]. In the standard central DP model, computing an approximate solution for the size of the maximum matching in the input graph with $O(\log(n)/\varepsilon)$ additive error is trivial since the sensitivity of the size of the maximum matching is $1$ and so we can use the Laplace mechanism to obtain such an estimate.
As discussed, the closest paper to ours is the private allocation paper of [hsu2014private]. They give a node-joint differentially private approximation algorithm that gives an approximate maximum matching solution on bipartite graphs. They model any bipartite graph as a set of bidders (on one side) and goods (on the other). Then, they formulate a differentially private version of the ascending auction price algorithms of [kelso1982job] to compute Walrasian equilibrium prices of the goods (a set of prices ensuring optimal utility of buyers and no good is over-demanded). They release differentially private counts of the number of bids each good has received using the private counting algorithms of [CSS11, DNPR10].
@@ 2.1 Differential Privacy
In particular, one can easily show that if every node processes the information contained in the public billboard using a deterministic algorithm, then the explicit solution contained at every node will be $\varepsilon$-edge DP (with respect to edge-neighboring graphs) except for the two nodes adjacent to the edge that differs between neighboring graphs. Morally speaking, such a set of explicit solutions does not leak any additional private information since the nodes that are endpoints to the edge that differs already know that this edge exists. Note that such a guarantee is stronger than joint differential privacy. Hence, we decouple the privacy definition from the solution release model and say an algorithm is $\varepsilon$-LEDP in the billboard model if the public billboard is produced via a $\varepsilon$-LEDP algorithm and the explicit solutions obtained by every node is via a deterministic algorithm at each node.
@@ 2.3 Differential Privacy Tools
In this section, we state the privacy tools we use in our paper. The adaptive Laplace mechanism is a formalization of the Laplace mechanism for adaptive inputs that we employ in this work (and is used implicitly in previous works). For completeness, we include a proof of Lemma 2.14 in Appendix A.
Lemma 2.14 (Adaptive Laplace Mechanism (used implicitly in [wagaman2024time])).
@@
Finally, we use the following privacy amplification theorem from subsampling.
Lemma 2.17 (Privacy Amplification via Subsampling Theorem).
If elements from the private dataset are sampled with probability $p$ and we are given a $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ on the original dataset, then running $\mathcal{A}$ on the subsampled dataset gives a $(2p\varepsilon,p\cdot\delta)$-DP algorithm for $\varepsilon\in(0,1)$.
2.4 Concentration Inequalities
Lemma 2.18.
Theorem 2.19 (Multiplicative Chernoff Bound).
  1. 1.
  2. 2.

### v2

@@ 1.1.4 Continual Release
We give results for two different types of streams in the continual release model. The first type of stream we consider is the arbitrary edge-order stream where the edges in the stream appear one by one in arbitrary order. In particular, in arbitrary edge-order streams, we show the following results for edge- and node-DP in Section 8. In our results below, we give bicriteria approximate solutions where an $(\psi,\phi)$-approximation is one with a $\psi$-multiplicative error and $\phi$-additive error. Our approximations hold, with high probability, for every update.
@@
Let $\eta\in(0,1]$ and $\varepsilon\in(0,1)$. Algorithm 6 is an $\varepsilon$-edge DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution that has the size of a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
@@
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{3}\varepsilon^{2}}\right)$and (ii) each implicit solution has the size of a
$\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta^{2}\varepsilon}\right)\right)$-approximate maximum matching.
The second type of stream we consider is the arbitrary adjacency-list order model [kallaugher2019complexity, goldreich2009proximity, czumaj2016relating, mcgregor2016better] in which nodes arrive in arbitrary order and once a node $v$ arrives as an update in the stream all edges adjacent to $v$ arrive in an arbitrary order as edge updates immediately after the arrival of $v$. In this model, we give the following results within Section 9.
@@
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, Algorithm 8 is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs, with high probability, an implicit $b$-matching that has the size of a $(2,1)$-approximate maximum matching.
@@
Model $b$ Approximation Bound Type Reference
$(\varepsilon,\delta)$-edge DP $O(1)$ $\Omega(n/b)$ Lower Bound Theorem 1.1
(explicit solutions)
$\varepsilon$-edge DP $\Omega(\log(n)/\varepsilon)$ $(2+\eta)$ Lower Bound Theorem 1.7
$\varepsilon$-LEDP ($n$ rounds) $O(\log(n)/\varepsilon)$ $2,(2+\eta)$ Upper Bound Theorem 1.3
Theorem 1.4
Theorem 1.5
$\varepsilon$-LEDP ($O(\log n)$ rounds) $O(\log^{2}(n)/\varepsilon)$ $2$ Upper Bound Theorem 1.6
$\varepsilon$-node DP $O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+{\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}}\right)$ $\left(2+\eta\right)$ Upper Bound Theorem 1.8
$\varepsilon$-node DP (bipartite) $s=\Omega\left(\frac{\log(n)}{{\eta^{4}}\varepsilon}\right)$ $(1+\eta)$, $s$-matching Upper Bound Theorem 1.10
Edge-Order $\varepsilon$-Edge DP $O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Theorem 1.11
@@
Edge-Order $\varepsilon$-Node DP $O\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+{\frac{\log^{3}(n)}{\eta^{3}\varepsilon^{2}}}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Theorem 1.12
@@
Adj-List Edge $O(\log(n)/\varepsilon)$ $(2,1)$ Upper Bound Theorem 1.13
@@
Table 1: Summary of results on differentially private matchings. All approximations are given in terms of the optimum maximum matching in the input graph. For our node-DP results, $\alpha$ denotes the arboricity of the graph. All results except the first row are with respect to implicit solutions while the first one is a lower bound for explicit solutions.
1.2 Technical Contributions of Independent Interest.
Our algorithms not only use ideas from parallel and distributed algorithms but also use several novel algorithmic tools, which may be of independent interest outside of the scope of this paper. In this section, we describe such novel contributions. In Section 1.3, we describe our contributions with respect to differentially private matchings.
Our lower bound proof relies on a symmetrization argument that enables us to prove the lower bound for a restricted class of highly symmetric algorithms and extend the hardness to all algorithms. To the best of our knowledge, this is the first application of symmetrization arguments towards DP lower bounds and may be of use outside of private matchings.
Our Public Vertex Subset Mechanism (Algorithm 4) allows us to privately release a set of candidate vertices per node. Each node can then intersect the released set with its private adjacency list to produce implicit vertex subset solutions. Our mechanism is an independent contribution that may be useful for other implicit graph problems beyond matching.
In order to implement our algorithms in the node-DP setting, we make novel use of matching sparsifiers for bounded-arboricity graphs [solomon2021local] by providing a “stable” implementation in the sense of [wagaman2024time] that allows us to apply our edge-DP algorithms in a black-box fashion on the sparsified input, while incurring a factor of the arboricity $\alpha$ in the approximation guarantees. While sparsification has been studied in the context of node privacy [wagaman2024time], previous utility guarantees typically depend on the maximum degree, which can be significantly bigger than the arboricity. [^5] The star graph on $n$ vertices has maximum degree $n-1$ but arboricity $\alpha=1$. We also prove a novel extension of the sparsifier to $b$-matchings, which may be of interest beyond private graph algorithms to the non-private graph algorithm setting. In general, bounded arboricity sparsifiers may be of use in node-DP beyond matchings.
@@ 1.3.1 Edge Privacy
Our main algorithms (Theorem 1.3 and Theorem 1.6) take inspiration from greedy algorithm and distributed algorithms for maximal matching, although our proofs also hold for $b$-matchings. In particular, the traditional greedy algorithm for maximal matching iterates through the nodes in the graph in an arbitrary order and matches the nodes greedily whenever there exists an unmatched node in the current iteration. However, this algorithm is clearly not private because each matched edge reveals the existence of a true edge in the graph. Thus, we must have some way of releasing public subsets of vertices that include both true edges and non-edges.
Instead, we introduce a novel DP proposal/receiver mechanism that allows nodes to propose a match set and receivers to accept the matches. We detail this proposal/receiver mechanism in Algorithm 2 and the guarantees are given in Theorem 1.3. Algorithm 2 proceeds in $n$ rounds using an arbitrary ordering of the nodes in the graph. We give as input a threshold $b$ which is a threshold for the number of edges in the matching each vertex is incident to. In the order provided by the ordering, each node $v$ privately checks using the Multidimensional AboveThreshold (MAT) technique introduced in [DLL23] whether it has reached its number of matched neighbors threshold. If it has not, then it asks all of its neighbors ordered after it whether they will match with it. This is the proposal phase of the algorithm. Each of these neighbors first determines using MAT whether they have exceeded their threshold for matching and then flips a coin (using progressively exponentially decreasing probability) to determine whether to be part of the set to match with $v$. Then, every vertex ordered after $v$ answers YES or NO depending on its coin flip. Intersecting the YES/NO responses with $v$’s private adjacency list results in $v$’s matches. This second portion of the algorithm is the receiver phase. All of these procedures can be done locally privately in the LEDP model.
@@
To speed up our algorithm to run in polylogarithmic rounds (Theorem 1.6), we need to allow many vertices to simultaneously propose matches to their neighbors (rather than just one at a time). Intuitively, this can be done in ways similar to classical parallel and distributed algorithms for related problems like Boruvka’s parallel MST algorithm [Boruvka]. In these algorithms, nodes randomly choose whether to propose or to listen. Then, proposer nodes send proposals, receivers decide which of their proposals to accept, and receivers communicate the acceptances back to the proposers (see [DHIN19] for a recent example of this in distributed settings). However, the added LEDP constraint makes this process significantly more difficult. Note, for example, that even communicating “with neighbors” violates privacy, since the transcript would then reveal who the neighbors are! Hence, greater care must be taken regarding the messages each node transmits. Fortunately, by using techniques gleaned from our $n$-round algorithm, ideas from basic parallel algorithms, and improved analyses, our implementation goes through.
The crux of our $O(\log n)$ round distributed algorithm relies on our Public Vertex Subset Mechanism. We use this mechanism to release proposal sets on the part of the proposers and match sets on the part of the receivers. The proposers release public proposal sets that contain proposals to its neighbors to match. Using the Public Vertex Subset Mechanism, receivers can privately discern which nodes have proposed. Then, receivers also publish a public subset of nodes as their match sets. Using this mechanism, the proposers can then (privately) recognize which receivers accepted their proposals. Performing multiple rounds of this procedure leads to our desired Theorem 1.6. Our novel utility analyses depends on a novel noisy version of a charging argument, based on edge orientation, implicitly described in [khuller1994primal].
@@ 1.3.2 Node Privacy
So if we could compute and release a differentially private matching sparsifier, we would be done. Unfortunately, we show in Section 11.2 that this is impossible, by providing a strong lower bound against private matching sparsification (Theorem 11.2). However, we show that we can use these sparsifiers in a node-DP algorithm even though releasing them does not satisfy DP! Intuitively, this is because given two node-neighboring graphs, computing a sparsifier of them can bring the distributions of the outputs “closer together”. In particular, although these sparsifiers cannot be released publicly, they decrease the “edge-distance” between node-neighboring graphs, where the edge-distance is the number of edges that differ between neighboring graphs.
We extend our results to $b$-matchings using our novel arboricity sparsifiers for $b$-matchings in Section 10.2.
Improvement in Bipartite Graphs. (Section 7)
The starting point of our algorithm for node-private bipartite matchings is the same deferred acceptance type algorithm which [hsu2014private] base their private algorithm on. In their private implementation of the algorithm, they use continual counters [dwork2009complexity, CSS11] to keep track of the number of bidders each item is matched to. This is necessary for their implementation since their implicit solutions need to know the number of bidders an item is matched with at every iteration of the algorithm. We show that in a different private implementation of the algorithm, the counts of the number of bidders an item is matched to is only needed $O(1/\eta^{2})$ different times (where $\eta\in(0,1]$ is present in the $(1+\eta)$-approximation factor). Furthermore, a clever use of MAT to iteratively check if this count exceeds the threshold $s$ (where $s$ is the matching factor) suffices for the implementation, enabling us to substantially improve the error of the counts used.
1.3.3 Continual Release (Sections 8 and 9)
@@ 1.4 Related Work
Private Allocations.
[…]
Now, auction-based methods and the implicit solution concept of [hsu2014private] are inherently tied to bipartite graphs. While it is possible to reduce general matchings to a bipartite instance via random bipartitions [liu2024pointwise], this transformation incurs a factor of 2 in the multiplicative approximation factor and the obtained approximation ratio is only in expectation. This would imply a node-DP $(2+\eta)$-approximate $s$-matching in expectation for general graphs. Our novel solution concept allows us to handle general graphs, under edge-DP or node-DP assuming low arboricity, and obtain a $(2+\eta)$-approximate $s$-matching with high probability while reducing the lower bound to $s=\Omega(\log(n)/\varepsilon)$. Moreover, specializing our techniques to the bipartite case (with arbitrary arboricity) recovers the node-DP $(1+\eta)$-approximation ratio of [hsu2014private], with the same reduced bound on $s$.
We remark that a followup work [hsu2016jointly] studies a class of private convex programming problems that capture fractional allocations with supply $s$ under approximate differential privacy. Similar to our algorithm, they also output an $s^{\prime}$-matching with $s^{\prime}>s$. Unlike our results, they compute a fractional allocation and guarantee that the sum of constraint violations ($\sum_{j}s_{j}^{\prime}-s_{j}$) is bounded, whereas our algorithm returns an (integral) solution while ensuring each constraint violation is bounded ($\max_{j}s_{j}^{\prime}-s_{j}$). Thus, neither the privacy nor the utility guarantees are comparable to ours.
Although [kannan2015pareto] also study private allocations, they do so in a relaxed notion of differential privacy called marginal differential privacy with the goal of computing approximate pareto optimal allocations. Thus both their privacy and utility guarantees differ from our setting.
Private Matching Size.
Differentially private estimations of the size of the maximum matching have been recently studied in various privacy models including the continual release model [FHO21, wagaman2024time, raskhodnikova2024fully, epasto2024sublinear] and the sublinear model [BGM22]. In the standard central DP model, computing an approximate solution for the size of the maximum matching in the input graph with $O(\log(n)/\varepsilon)$ additive error is trivial since the sensitivity of the size of the maximum matching is $1$ and so we can use the Laplace mechanism to obtain such an estimate.
Implicit Solutions.
@@ 2.1 Differential Privacy
In particular, one can easily show that if every node processes the information contained in the public billboard using a deterministic algorithm, then the explicit solution contained at every node will be $\varepsilon$-edge DP (with respect to edge-neighboring graphs) except for the two nodes adjacent to the edge that differs between neighboring graphs. Morally speaking, such a set of explicit solutions does not leak any additional private information since the nodes that are endpoints to the edge that differs already know that this edge exists. Hence, we decouple the privacy definition from the solution release model and say an algorithm is $\varepsilon$-LEDP in the billboard model if the public billboard is produced via a $\varepsilon$-LEDP algorithm and the explicit solutions obtained by every node is via a deterministic algorithm at each node.
@@ 2.2 Differential Privacy Tools
In this section, we state the privacy tools we use in our paper. The adaptive Laplace mechanism is a formalization of the Laplace mechanism for adaptive inputs that we employ in this work (and is used implicitly in previous works). For completeness, we include a proof of Lemma 2.14 in Section A.2.
Lemma 2.14 (Adaptive Laplace Mechanism; used implicitly in [wagaman2024time]).
@@
Finally, we use the following privacy amplification theorem for subsampling.
Lemma 2.17 (Privacy Amplification via Subsampling).
Let $\varepsilon\in(0,1]$. If elements from the private dataset are sampled with probability $p$ and we are given a $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ on the original dataset, then running $\mathcal{A}$ on the subsampled dataset gives a $(2p\varepsilon,p\cdot\delta)$-DP algorithm for $\varepsilon\in(0,1)$.